OpenStudy (taylor):

i need help on 7x+y=-4... can anyone help please?... :D

OpenStudy (bahrom7893):

i can so whats the question?

OpenStudy (anonymous):

hello

OpenStudy (taylor):

im solving for y

OpenStudy (taylor):

and hello

OpenStudy (anonymous):

i got it

OpenStudy (bahrom7893):

y = -4-7x

OpenStudy (bahrom7893):

hi, can u please click on become a fan if i helped, thanks!

OpenStudy (anonymous):

In mathematics, the solution to the equation 7x -y + 4 is y = 7 x+4 where "x" and "y" are considered variables.

OpenStudy (taylor):

haha ok welll can you explain it where i i can understand cause im 14 and suck at math

OpenStudy (bahrom7893):

dkglenn wrong!

OpenStudy (anonymous):

what no i am not How

OpenStudy (bahrom7893):

7x+y=-4; y = -4-7x

OpenStudy (bahrom7893):

please click on become a fan taylor! Thank you!

OpenStudy (anonymous):

wrong

OpenStudy (taylor):

lol so who do i belive..ahaha

OpenStudy (bahrom7893):

dkglenn stop confusing people!! 7x+y=-4; subtract 7x from both sides y = -4-7x

OpenStudy (taylor):

lol its ok you boh were very helpful no biggy

OpenStudy (anonymous):

ok lol

OpenStudy (taylor):

thanks dkglenn

OpenStudy (taylor):

:D

OpenStudy (taylor):

i dont understand why i suck at these so much.. :/

OpenStudy (anonymous):

i do to lol! your not alone

OpenStudy (taylor):

lol im probably worse i have been taking this class for like two years and im sux and im cramiing for a test